OpenStudy (anonymous):

If and x = -12, find y.

OpenStudy (diyadiya):

There is something wrong with your question

OpenStudy (diyadiya):

Is there any equation?

OpenStudy (anonymous):

no If x = -12, find y.

OpenStudy (anonymous):

question is not complete

OpenStudy (diyadiya):

y can be anything lol check your question again!

OpenStudy (anonymous):

if y=1/4 and x=-12,find y

OpenStudy (diyadiya):

y is already given then why are you finding y again?

OpenStudy (anonymous):

if y=1/4x and x=-12,find y lol

OpenStudy (diyadiya):

Ahh! Now the question makes sense :D \[y = \frac{1}{4x} \]Plug in the value of x

OpenStudy (diyadiya):

like this \[y= \frac{1}{4 \times -12}\]

OpenStudy (anonymous):

thanks lol :)

OpenStudy (diyadiya):

got the answer?

OpenStudy (anonymous):

no

OpenStudy (diyadiya):

lol where are you stuck ? what is 4 x -12 ?

OpenStudy (anonymous):

how do you solve it with the -12

OpenStudy (anonymous):

-3

OpenStudy (diyadiya):

ok when you multiply two negative numbers the answer is positive when you multiply one positive number with a negative number the answer is negative for example 3 x -4 =-12 -3 x -3 = 9

OpenStudy (anonymous):

so it will be -3

OpenStudy (diyadiya):

How?

OpenStudy (diyadiya):

wait your question is \[y= \frac{1}{4}x\] or \[y= \frac{1}{4x}\]

OpenStudy (anonymous):

y=14x

OpenStudy (diyadiya):

y=14x? so y=14 x -12 can you do 14 x -12 ?

OpenStudy (diyadiya):

i've explained how

OpenStudy (diyadiya):

@hortokam

OpenStudy (anonymous):

The relationship between two quantities that increase or decrease together is called ______.

OpenStudy (diyadiya):

Well ,you didn't answer the previous one :D

OpenStudy (anonymous):

i did -12/4=-3

OpenStudy (diyadiya):

your question is \[y=\frac{x}{4}\]??????

OpenStudy (anonymous):

i got a another one Find the equation of (3,1)